# Relational Self-supervised Distillation with Compact Descriptors for Image Copy Detection

### v1

@@ 3.3. Contrastive Learning for ICD
3.4. The Hard Negative Loss
In the copy detection scenario, the primary challenges for training arise from the hard negative pairs and the fact that the use of small descriptors renders the environment susceptible to dimensional collapse. To address this, we employ the Hard Negative (HN) loss (Kordopatis-Zilos et al., 2023) to finely control the entropy and minimize the hard negative pairs. We also introduce an additional term that ensures that a hard negative pair can push their descriptors apart.
(7) $\mathcal{L}_{hn}=-\frac{1}{N}\log\sum_{i=1}^{N}max_{j\in n_{i,j}}(1-S_{i,j})$
@@ 3.4. The Hard Negative Loss
(8) $\mathcal{L}_{RDCD}=\lambda_{rel}\mathcal{L}_{rel}+\lambda_{con}\mathcal{L}_{con}+\lambda_{hn}\mathcal{L}_{hn}$
4. Experiments
4.1. Dataset
@@
DISC2021 is a dataset consisting of training, reference and query images, that is used for the Image Similarity Challenge (Douze et al., 2021b). The query image set contains strong auto-augmented and human-augmented images, and distractive images that may not appear in the reference set. The training set is designed to train a model without any labels, and the reference set is used for searching the background as a database.
@@
Copydays (Jegou et al., 2008) is a dataset designed to detect copied content, that consists of 157 original images and 2,000 queries with cropping, jpeg compressing and strong augmentations. We add 10k distractors from YFCC100M (Thomee et al., 2016), a common practice (Caron et al., 2021; Berman et al., 2019) that is known as CD10K. We evaluate our method based on the mean average precision(mAP) and micro average precision($\mu AP$) for the strongly transformed copies.
@@
NDEC (Wang et al., 2023) dataset incorporates HN distractors, making it more advanced than conventional ICD datasets. The basic data is derived from DISC2021 while the HN images are sourced from OpenImage. Similar to DISC2021, the NDEC dataset is organized into training, query, and reference sets. The training data includes 900,000 basic images and adds HN pairs (100,000 x 2 paired images) to enhance the difficulty level of ICD. The test data contains 49,252 query images and 1,000,000 reference images, with 24,252 of the query images being HN.
4.2. Evaluation Metrics
@@
To enable a fair comparison with state-of-the-art methods (Pizzi et al., 2022; Caron et al., 2021), we use $\mu AP$ (Perronnin et al., 2009), accuracy at 1, and recall at precision 90 as the evaluation metrics. The equation for $\mu AP$ is as follows:
(9) $\mu AP=\sum\limits_{i=1}^{N}P(i)\triangle r(i)\in[0,1]$
@@
where $P(i)$ is the precision at position $i$ of the sorted list of pairs, $\triangle r(i)$ is the difference in the recall between position i and i-1, and N is the total number of returned pairs for all queries. Any detected pair for a distractor query will decrease the average precision, thus all queries are evaluated together by merging the returned pairs for all queries, sorting them by confidence, and generating a single precision-recall curve. This is different from mAP, also known as macro-AP (Perronnin et al., 2009), where the average precision is computed separately per query and then averaged over all queries. $\mu AP$ considers confidence values to capture all queries, while mAP only considers the query in the ground truth. Thus, $\mu AP$ is a more accurate metric in our context because our queries contain the images not contained in the references. The former is more appropriate for ICD, while the latter is typically used in retrieval tasks.
[…]
Method Network Size $\mu$AP $\mu$APSN
Multigrain RN-50 1500 16.5 36.5
Multigrain RN-50 2048 20.5 41.7
DINO ViT-B/8 1536 32.6 54.4
DINO ViT-B/16 1536 32.2 53.8
SSCD RN-50 512 61.5 72.5
SSCD ResNeXt-101 1024 63.7 75.3
@@ 4.3. Training Implementation
We train all networks using the pretraining parameters from ImageNet-1K. Batch normalization statistics are synchronized across all GPUs. We adhere to the hyperparameters outlined in SimCLR (Chen et al., 2020b), MoCo-v2 (Chen et al., 2020a) and SSCD (Pizzi et al., 2022). The batch size is set at N = 256, with a resolution of 224 x 224 for training and 288 x 288 for inference. We use a learning rate of 1.0 and a weight decay of $10^{-6}$. We train all models for 100 epochs. We employ a cosine learning rate scheduler with a warmup period of five epochs and use the Adam optimizer. For our knowledge distillation process, we set the temperature at 0.04 for the teacher and 0.07 for the student. All experiments are conducted on an NVIDIA A100 80GB GPU. To ensure reproducibility, we maintain consistent seed values throughout all of the experiments.
4.4. Results
Table 1 presents the results for the DISC2021 dataset. The first section of the table summarizes the performance of different methods with varying descriptor sizes, while the second section describes the performance of our baseline SSCD (Pizzi et al., 2022) strategy using a lightweight network with descriptor sizes of 64, 128, and 256. When training using a lightweight network, we do not use a mix-up for the fair comparison of our methods. The third section shows the results of the proposed RDCD approach.
When compared with the methods outlined in the first section, our RDCD method achieves competitive performance despite using significantly smaller descriptors and a more compact network size. It is noteworthy that the RDCD method, with a descriptor size of 64, achieves a $\mu$APSN of 53.5, which is comparable to the DINO method that utilizes a ViT-B/16 network with a descriptor size of 1536, yielding a $\mu$APSN of 53.8. This comparison is significant, as the descriptor size of RDCD is 24 times smaller.
@@ 4.4. Results
When evaluated against an equivalent lightweight architecture, RDCD consistently outperforms SSCD across all three dimensions assessed. We employ the SSCD model RN-50, which has a $\mu$APSN of 72.5, as the teacher network to provide knowledge to the lightweight network. When we apply our methods to EfficientNet-B0, RDCD outperforms SSCD. Remarkably, RDCD with a descriptor size of 128 achieves a $\mu$APSN of 61.1, which is competitive and matches the performance of SSCD with a descriptor size of 512. This indicates that our method is capable of delivering strong results even with smaller descriptors.
@@
RDCD also outperforms SSCD in terms of mAP using the CD10K dataset (Table 2). Specifically, when using EfficientNet-B0 with a descriptor size of 128, RDCD achieves an mAP of 79.2, which is significantly higher than the best result for SSCD. With a descriptor size of 256, the performance improvement is even more pronounced, with RDCD reaching a mAP of 81.4.
@@
For the NDEC dataset, Table 3 shows that our method shows competitive performance compared to the teacher network SSCD-RN50, which achieves a $\mu$APSN of 23.3 with a descriptor size of 512. Our proposed RDCD approach, employing EN-B0 network with descriptor sizes of 64, 128, and 256 achieves a $\mu$APSN of 17.3, 19.5, and 21.3, respectively. This demonstrates that across all compact descriptor sizes, RDCD surpasses the performance of the previous SSCD when using the same lightweight architecture.
@@
Table 2. Results on the CD10K (Copydays+ 10k distractors). We compare different methods, network architectures (ResNet-50, ViT-B/8, ViT-B/16, ResNeXt-101, EfficientNet-B0), model sizes, and performance metrics (mean Average Precision (mAP) and micro Average Precision ($\mu$AP). * means our implementation.
Method Network Size mAP $\mu$AP
Multigrain RN-50 1500 82.3 77.3
DINO ViT-B/8 1536 85.3 91.7
DINO ViT-B/16 1536 80.7 88.7
SSCD RN-50 512 85.0 97.9
SSCD ResNext-101 1024 91.9 96.5
using lightweight
@@
Table 3. Comparison of NDEC results. * means our implementation.
Method Network Size $\mu$AP $\mu$APSN
DINO ViT-B/8 1536 0.08 11.4
DINO ViT-B/16 1536 0.09 13.1
SSCD RN-50 512 21.2 23.3
@@
SSCD EN-B0 64 14.4 16.4
SSCD EN-B0 128 16.5 18.7
SSCD EN-B0 256 17.5 20.1
SSCD EN-B0 512 17.3 20.4
@@
RDCD EN-B0 64 15.7 17.3
RDCD EN-B0 128 17.5 19.5
RDCD EN-B0 256 18.8 21.3
5. Discussion
5.1. Effect of RSD and HN Loss
Our method employs an additional projector to reduce the descriptor size, which creates a network with a two-layer MLP. This configuration can lead to implicit regularization, due to the interaction between the weight matrices across different layers. Implicit regularization typically restricts the ability of a network to learn diverse features, leading to dimensional collapse with contrastive SSL (Jing et al., 2021). However, when applying HN loss together with contrastive learning, we show that dimensional collapse does not occur even with the use of an MLP, thus reducing the impact of implicit regularization.
Figure 3. Log of singular values for a descriptor size of 256, with and without HN loss.
To verify this, we calculate the singular values of the final descriptor of a model with a descriptor size of 256 under four conditions: with the use of RSD, FKD and with and without the use of HN loss. We compute the descriptors for 50k queries from the DISC21 dataset, calculate their covariance, and perform singular value decomposition on the covariance matrix. Subsequently, we extract the singular values and apply a logarithmic operation to them. As illustrated in Figure 3, RDCD has a full rank, while the absence of HN loss results in dimensional collapse. It can also be observed that applying HN loss with FKD does not have a significant effect.
During the training, we visualize the similarity of the hard negative for both the MoCo-v2 baseline and HN loss, implemented with either RSD or FKD, using descriptor sizes of 128 or 256, as shown in Figure 4. FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error (MSE) loss. As Figure 4 shows, using FKD with HN loss fails to reduce the similarity of the nearest negatives, whereas using RSD with HN loss gradually reduces the nearest negative similarity.
We further visualize the difference between positive similarity and hard negative similarity across multiple dimensions experimented in MoCo-v2 — RSD environment, comparing scenarios with and without the application of HN loss. We analyze a sample of 5,000 distinct queries extracted from the DISC21 dataset, which are separate from the queries used in the standard evaluation process for generality, to calculate and assess the differences in the similarity between positive and hard negative samples. As Figure 5 shows, across all descriptor sizes, the application of HN loss resulted in a more pronounced disparity compared to when it was omitted. This observation implies that HN loss effectively enhances the separation between positive and hard negative samples. Furthermore, we noted that as the dimensionality increases, the difference between positive and hard negative similarity also enlarges. We conjecture that this trend may contribute to improved performance in higher-dimensional descriptors.
Figure 4. Comparison of hard negative similarity of RSD and FKD.
@@ 5.1. Effect of RSD and HN Loss
Figure 5. Comparison of the difference in similarity between positives and nearest negatives with and without the use of HN Loss.
Table 4. Comparison of other distillation. All methods are trained with EfficientNet-B0 network. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
128 descriptors
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD 43.5 56.2 - -
SimCLR 46.2 58.7 48.4 59.9
47.4 59.6 49.6 61.4
47.5 59.6 50.8 61.5
MoCo-v2 40.3 57.5 44.8 59.1
47.1 60.7 50.0 61.1
46.2 59.3 49.6 61.0
256 descriptors
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD 46.0 59.8 - -
SimCLR 51.7 64.0 51.4 64.5
52.2 65.6 52.7 65.8
52.4 65.5 49.9 64.6
MoCo-v2 41.3 56.6 46.1 60.1
47.2 61.0 52.9 65.7
47.3 60.8 53.2 65.9
5.2. Comparison with Other Forms of Distillation
To further demonstrate the efficacy of RSD, we conduct a comparative analysis with FKD by 1) employing FKD alone, 2) employing RSD alone, and 3) integrating both FKD and RSD (Table 13). We also employ SimCLR-style contrastive learning. When FKD and RSD are combined, the best performance is observed in several cases, suggesting that combining both distillation methods can further improve performance. Nevertheless, RSD consistently outperforms FKD in all cases, with RSD alone achieving performance on par with the combined use of both distillation methods. This indicates that RSD is sufficiently robust, and FKD does not significantly alter the outcome. It confirms that RSD can operate alone to enhance the performance of compact descriptors. Additional experiments with other distillation methods are presented in Appendix A.
5.3. Effect of Different Architectures
We demonstrate the effectiveness of our proposed approach in scenarios where the teacher networks belong to different architectural families (ViT or CNN), with student models also selected from these two distinct architectures. For the ViT, we use DINO as the teacher, and for the CNNs, we use SSCD-RN50 and SSCD-ResNeXt-101. All models are trained on SSCD-RN50 hyperparameter settings. With DINO as the teacher model, we follow the ICD settings presented in (Caron et al., 2021) for the DINO baseline. Consequently, we distill a descriptor size of 1536 from the teacher (concatenation of 768 class tokens and 768 GeM pooled patch tokens). In the case of ResNeXt-101 teacher model, we distill the final target descriptor with a size of 1024. Table 5 shows that our method is highly effective for the different architectures, highlighting its adaptability and robustness.
Table 5. DISC evaluation with different architectures. All experiments are performed with a descriptor size of 128.
T S Eff-B0 FastViT-T12 Mob-V3
$\mu$AP $\mu$APSN $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
DINO-ViT-B/8 32.1 52.1 28.1 52.2 39.6 54.7
SSCD-RN-50 50.0 61.1 48.5 61.1 47.3 58.8
SSCD-ResNeXt-101 44.9 59.9 42.8 57.4 41.1 56.6
5.4. Ablation Study

### v5

@@ 3.3 Contrastive Learning for ICD
3.4 The Hard Negative Loss
In the copy detection scenario, the primary challenges for training arise from the hard negative pairs and the fact that the use of small descriptors renders the environment susceptible to dimensional collapse. To address this, we employ the Hard Negative (HN) loss [21] to finely control the entropy and minimize the hard negative pairs. We also introduce an additional term that ensures that a hard negative pair can push their descriptors apart.
$\mathcal{L}_{hn}=-\frac{1}{N}\log\sum_{i=1}^{N}max_{j\in n_{i,j}}(1-S_{i,j})$ (7)
@@ 3.4 The Hard Negative Loss
$\mathcal{L}_{RDCD}=\lambda_{rel}\mathcal{L}_{rel}+\lambda_{con}\mathcal{L}_{con}+\lambda_{hn}\mathcal{L}_{hn}$ (8)
4 Experiments
4.1 Dataset
@@
DISC2021 is a dataset consisting of training, reference and query images, that is used for the Image Similarity Challenge [8]. The query image set contains strong auto-augmented and human-augmented images, and distractive images that may not appear in the reference set. The training set is designed to train a model without any labels, and the reference set is used for searching the background as a database.
@@
Copydays [17] is a dataset designed to detect copied content, that consists of 157 original images and 2,000 queries with cropping, jpeg compressing and strong augmentations. We add 10k distractors from YFCC100M [40], a common practice [5, 3] known as CD10K. We evaluate our method based on the mean average precision(mAP) and micro average precision($\mu AP$) for the strongly transformed copies.
@@
NDEC [44] dataset incorporates HN distractors, making it more advanced than conventional ICD datasets. The basic data is derived from DISC2021 while the HN images are sourced from OpenImage. Similar to DISC2021, the NDEC dataset is organized into training, query, and reference sets. The training data includes 900,000 basic images and adds HN pairs (100,000 x 2 paired images) to enhance the difficulty level of ICD. The test data contains 49,252 query images and 1,000,000 reference images, with 24,252 of the query images being HN.
4.2 Evaluation Metrics
@@
To enable a fair comparison with state-of-the-art methods [31, 5], we use $\mu AP$ [30] as the evaluation metric. The equation for $\mu AP$ is as follows:
$\mu AP=\sum\limits_{i=1}^{N}P(i)\triangle r(i)\in[0,1]$ (9)
@@
where $P(i)$ is the precision at position $i$ of the sorted list of pairs, $\triangle r(i)$ is the difference in the recall between position i and i-1, and N is the total number of returned pairs for all queries. Any detected pair for a distractor query will decrease the average precision, thus all queries are evaluated together by merging the returned pairs for all queries, sorting them by confidence, and generating a single precision-recall curve. This is different from mAP, also known as macro-AP [30], where the average precision is computed separately per query and then averaged over all queries. $\mu AP$ considers confidence values to capture all queries, while mAP only considers the query in the ground truth. Thus, $\mu AP$ is a more accurate metric in our context because our queries contain the images not contained in the references. The former is more appropriate for ICD, while the latter is typically used in retrieval tasks. We additionally evaluate with $\mu AP_{SN}$. Score normalization is a post-processing technique commonly used in image search to achieve a more uniform distribution. It adjusts the scores based on an estimate of the density around the query and normalizes them using background images that were not involved in training.
Table 1: Results for the DISC2021 showing different methods, network architectures (RN-50, ViT-B/8, ViT-B/16, ResNeXt-101, EfficientNet-B0), model sizes, and performance metrics (micro Average Precision $\mu$AP and micro Average Precision with score normalization $\mu$APSN). * means our implementation. $+$ used as teacher model in our experiments.
Method Network $dim$ $\mu$AP $\mu$APSN
Multigrain [3] RN-50 1500 16.5 36.5
Multigrain [3] RN-50 2048 20.5 41.7
DINO+ [5] ViT-B/8 1536 32.6 54.4
DINO [5] ViT-B/16 1536 32.2 53.8
SSCD+ [31] RN-50 512 61.5 72.5
SSCD+ [31] ResNeXt-101 1024 63.7 75.3
@@ 4.3 Training Implementation
We train all networks using the pretraining parameters from ImageNet-1K. Batch normalization statistics are synchronized across all GPUs. We adhere to the hyperparameters outlined in SimCLR [6], MoCo-v2 [7] and SSCD [31]. The batch size is set at N = 256, with a resolution of 224 x 224 for training and 288 x 288 for evaluation. We use a learning rate of 1.0 and a weight decay of $10^{-6}$. We train all models for 100 epochs. We employ a cosine learning rate scheduler with a warmup period of five epochs and use the Adam optimizer. For our knowledge distillation process, we set the temperature at 0.04 for the teacher and 0.07 for the student. All experiments are conducted on 1 NVIDIA A100 80GB GPU. To ensure reproducibility, we maintain consistent seed values throughout all of the experiments. We further discuss the practicality on RDCD in Appendix E.
4.4 Results
@@
Table 1 presents the results for the DISC2021 dataset. The first section of the table summarizes the performance of different methods with varying descriptor sizes, while the second section presents the performance of our baseline SSCD [31] strategy using a lightweight network with descriptor sizes of 64, 128, and 256. When training using a lightweight network, we do not use mix-up to ensure a fair comparison of our methods. The third section shows the results of the proposed RDCD approach. When compared with the methods outlined in the first section, our RDCD method achieves competitive performance despite using significantly smaller descriptors and a more compact network size. It is noteworthy that the RDCD method, with a descriptor size of 64, achieves a $\mu$APSN of 53.5, which is comparable to the DINO method, which utilizes a ViT-B/16 network with a descriptor size of 1536, yielding a $\mu$APSN of 53.8. This comparison is significant, as the descriptor size of RDCD is 24 times smaller. When evaluated with an equivalent lightweight architecture, RDCD consistently outperforms SSCD across all three dimensions assessed. We use the SSCD model RN-50, which has a $\mu$APSN of 72.5, as the teacher network to transfer knowledge to the lightweight network. When we apply our methods to EfficientNet-B0, RDCD outperforms SSCD. Remarkably, RDCD with a descriptor size of 128 achieves a $\mu$APSN of 61.1, which is competitive and matches the performance of SSCD with a descriptor size of 512. This indicates that our method is capable of delivering strong results even with smaller descriptors.
@@
RDCD also outperforms SSCD using the CD10K dataset in Table 2. Specifically, when using EfficientNet-B0 with a descriptor size of 128, RDCD achieves an mAP of 79.2, which is significantly higher than the best result for SSCD. With a descriptor size of 256, the performance improvement becomes more significant, achieving an mAP of 81.4.
@@
For the NDEC dataset, Table 3 shows that our method shows competitive performance compared to the teacher network SSCD-RN50, which achieves a $\mu$APSN of 23.3 with a descriptor size of 512. Our proposed RDCD approach, employing EN-B0 network with a descriptor size of 64, 128, and 256 achieves a $\mu$APSN of 17.3, 19.5, and 21.3, respectively. This demonstrates that across all compact descriptor sizes, RDCD surpasses the performance of the previous SSCD when using the same lightweight architecture.
@@
Table 2: Results on the CD10K (Copydays+ 10k distractors). We compare different methods, network architectures (RN-50, ViT-B/8, ViT-B/16, ResNeXt-101, EfficientNet-B0), model sizes, and performance metrics (mean Average Precision (mAP) and micro Average Precision ($\mu$AP). * means our implementation.
Method Network $dim$ mAP $\mu$AP
Multigrain [3] RN-50 1500 82.3 77.3
DINO [5] ViT-B/8 1536 85.3 91.7
DINO [5] ViT-B/16 1536 80.7 88.7
SSCD [31] RN-50 512 85.0 97.9
SSCD [31] ResNeXt-101 1024 91.9 96.5
@@
Table 3: Comparison of NDEC results. * means our implementation.
Method Network $dim$ $\mu$AP $\mu$APSN
DINO [5] ViT-B/8 1536 16.2 22.8
DINO [5] ViT-B/16 1536 18.1 26.2
SSCD [31] RN-50 512 42.4 46.6
@@
SSCD EN-B0 64 28.9 32.8
SSCD EN-B0 128 33.0 37.5
SSCD EN-B0 256 35.0 40.1
SSCD EN-B0 512 34.5 40.8
@@
RDCD EN-B0 64 31.4 34.7
RDCD EN-B0 128 34.9 38.9
RDCD EN-B0 256 37.5 42.5
5 Discussion
5.1 Effect of RSD with HN Loss
Our method employs an additional projector to reduce the descriptor size, which creates a network with a two-layer MLP. This configuration can lead to implicit regularization, due to the interaction between the weight matrices across different layers. Implicit regularization typically restricts the ability of a network to learn diverse features, leading to dimensional collapse with contrastive SSL [18]. However, when applying HN loss together with RSD, we show that dimensional collapse does not occur even with the use of an MLP, thus reducing the impact of implicit regularization.
Figure 3: Log of singular values for a descriptor size of 256, with and without HN loss.
To verify this, we calculate the singular values of the final descriptor of a model with a descriptor size of 256 under four conditions: with the use of RSD, FKD and with and without the use of HN loss. We compute the descriptors for 50k queries from the DISC2021 dataset, calculate their covariance, and perform singular value decomposition on the covariance matrix. Subsequently, we extract the singular values and apply a logarithmic operation to them. As illustrated in Figure 3, RDCD has a full rank, while the absence of HN loss results in dimensional collapse. It can also be observed that applying HN loss with FKD does not have a significant effect. We further visualize the difference between positive similarity and hard negative similarity across multiple dimensions experimented in MoCo-v2 |||| RSD environment, comparing scenarios with and without the application of HN loss. We analyze a sample of 5,000 distinct queries extracted from the DISC2021 dataset, which are separate from the queries used in the standard evaluation process for generality, to calculate and assess the differences in the similarity between positive and hard negative samples. As Figure 4 shows, across all descriptor sizes, the application of HN loss resulted in a more pronounced disparity compared to when it was omitted. This observation implies that HN loss effectively enhances the separation between positive and hard negative samples. Furthermore, we noted that as the dimensionality increases, the difference between positive and hard negative similarity also enlarges. We conjecture that this trend may contribute to improved performance in higher-dimensional descriptors.
@@ 5.1 Effect of RSD with HN Loss
Figure 4: Comparison of the difference in similarity between positives and nearest negatives with and without the use of HN Loss.
5.2 Effect of Hard Negative Loss on Uniformity
A uniform distribution of descriptors represents semantically similar images far apart while placing copies of the same image close together. Both DINO [5] and SSCD [31] are used as teacher models; however, only SSCD uses entropy regularizer to achieve a uniform distribution. In general, features used during evaluation are called descriptors. We demonstrate that applying HN loss to the final feature used as a descriptor results in a more uniform distribution. This effect occurs regardless of whether the teacher was trained with or without HN loss. Figure 5 visually supports this observation. b) shows the distribution of a student model distilled from a non-uniform teacher and trained with HN loss, while c) displays the distribution of a student model distilled from a uniform teacher, also trained with HN loss. Both distributions are notably more uniform compared to the distribution of the DINO teacher in a). We further analyze the intermediate feature, which is directly distilled by the teacher model, and show that when using RSD, the uniformity of the teacher’s embedding is transferred to the student’s embedding. d) shows the result of distillation from a non-uniform teacher DINO, while e) shows the result of distillation from a uniform teacher SSCD. Consequently, if the teacher’s embedding has a uniform distribution in the embedding space, applying HN loss to the final feature may yield limited additional uniformity. Additional experiments with intermediate features are illustrated in Appendix B.
Figure 5: T-SNE visualization of descriptors from different model configurations: a) DINO baseline, b) DINO-MoCo-v2 w/ HN loss, c) SSCD-MoCo-v2 w/ HN loss. d) DINO-MoCo-v2 w/o HN loss, e) SSCD-MoCo-v2 w/o HN loss. In a),b) and c), the final features are used as descriptors whereas in d) and e), intermediate features are used. We randomly select 9 classes and visualize 50 images from each class from ImageNet.
5.3 Ablation Study